# Identifying and Mitigating API Misuse in Large Language Models

### v1

@@ 4.1 RQ1: How well do LLMs handle API invocation in code generation?
TABLE I: Performance comparison of StarCoder and Copilot in Python and Java for method and parameter prediction tasks.
@@
BLEU CodeBERTScore EM BLEU CodeBERTScore EM
Method StarCoder 84.6 92.3 2,268 81.2 86.4 1,328
Copilot 88.3 92.6 1,911 86.1 87.7 1,593
Parameters StarCoder 85.0 92.8 2,312 82.0 87.0 1,330
Copilot 89.0 93.0 1,956 87.0 88.2 1,617
@@
We first wanted to investigate how well our two studied LLMs can generate code to correctly invoke APIs in code generation. As described above, we sample 3,000 distinct code snippets in Python and Java, respectively, from The Stack dataset. After prompting StarCoder and Copilot to complete Parameter Prediction and Parameter Generation independently, we collect 24,000 samples in total. We measure the correctness of the API invocation using three metrics introduced in subsection 3.6: BLUE, CodeBERTScre, and EM. Table I shows the performance of StarCoder and Copilot on in-the-wild API usage tasks in Python and Java. The results reveal that Copilot consistently outperforms StarCoder across all metrics, with the EM metric providing the clearest indication of model capability in generating syntactically correct and precise code.
For Python, Copilot achieves an EM of 1911, compared to StarCoder’s 2268. This significant difference underscores Copilot’s superior ability to produce exact matches to reference implementations. Additionally, Copilot surpasses StarCoder in BLEU and CodeBERTScore, achieving scores of 88.3 and 92.6, respectively, compared to StarCoder’s 84.6 BLEU and 92.3 CodeBERTScore. These results highlight that Copilot is more effective in generating both syntactically accurate and semantically similar Python code. The higher EM score further solidifies its advantage in tasks requiring exact code generation.
In the Java domain, the trend remains consistent, with Copilot outperforming StarCoder in all metrics. For Java, Copilot achieves an EM of 1593, compared to StarCoder’s 1328, again reflecting its ability to generate more precise API usage code. Similarly, Copilot’s BLEU score of 86.1 and CodeBERTScore of 87.7 surpass StarCoder’s BLEU of 81.2 and CodeBERTScore of 86.4. These results suggest that Copilot is better suited for Java tasks, producing fluent and semantically aligned code with higher precision.
The comparison between these two decoder-based models highlights the significance of pretraining strategies and dataset scale. Copilot’s consistently higher EM scores across both Python and Java suggest it is better optimized for tasks that demand exact code matches, while its performance in BLEU and CodeBERTScore indicates strong fluency and semantic alignment. Although StarCoder performs competitively in CodeBERTScore, its lower EM scores in both languages suggest room for improvement in generating syntactically exact outputs.
437StarCoderCopilot1721235
(a) Wrong Method Inference in Java
190StarCoderCopilot547542
(b) Wrong Method Inference in Python
665StarCoderCopilot3781005
(c) Wrong Parameter Generation in Java
237StarCoderCopilot593451
(d) Wrong Parameter Generation in Python
@@
Figure 8: Comparison of incorrect API usage (wrong method inference and parameter generation) between StarCoder and Copilot in Java and Python.
Figure 8 shows the overlap and discrepancies in API misuse between StarCoder and Copilot. These Venn diagrams illustrate instances where one model produces incorrect API usage, either in the method call or parameter context, for Python and Java. In Python, Copilot demonstrates a clear advantage over StarCoder in API method inference, as shown in 8b. StarCoder has 190 unique errors, compared to Copilot’s 547, with 542 errors shared between the two models. These shared errors highlight common challenges, likely stemming from ambiguous requirements or limitations in the training datasets. (8a) shows a similar trend for Java, where StarCoder produces 437 unique method errors versus Copilot’s 172, with 1235 shared errors. This suggests that both models face consistent difficulties in method inference due to the complexity of Java APIs. For parameter generation, the patterns differ between languages. In Python (8d), StarCoder has 237 unique parameter errors, while Copilot has 593, with 451 shared errors. In Java (8c), StarCoder generates 665 unique parameter errors compared to Copilot’s 378, with 1005 shared errors, indicating distinct challenges in precise parameter usage across both languages.
@@
Finding 1: While Copilot demonstrates higher BLEU and CodeBERTScore metrics across both languages, error patterns vary significantly by language. Copilot shows fewer API misuses than StarCoder in Java but makes more errors in Python for both method inference and parameter generation. Both models share a substantial number of common errors, indicating persistent challenges in handling complex API usage scenarios.
@@ 4.2 RQ2: How do LLMs misuse APIs in the wild?
TABLE II: Comparison of API Misuse by StarCoder in Java and Python. “None” misuse type means that the LLM-based API usage is considered appropriate based on human inspection.
@@
TABLE III: Comparison of API Misuse by Copilot in Java and Python. “None” misuse type means that the LLM-based API usage is considered appropriate based on human inspection.
@@
Acceptable 35.0 24.5 28.0 27.3
We further analyze the patterns and distributions of API misuse of LLMs. Specically, we study the annotated outputs from both StarCoder and Copilot in subsection 3.7, which failed on EM score. Table II and Table III present a detailed analysis of how different types of API misuse manifest in Java and Python code generated by StarCoder and Copilot, respectively. The tables categorize the misuses into several types and compare the percentage occurrences of each type for both programming languages.
In both Java and Python, the most common misuse type is Hallucination, with significant differences in frequency between the two languages. For StarCoder, hallucinations account for 43.3% of parameter misuses and 34.6% of method misuses in Java, while for Python, the hallucination rate is even higher: 31.3% for parameters and 40.9% for methods. For Copilot, hallucinations are slightly lower but still prominent, with 39.0% of parameter misuses and 29.0% of method misuses in Java, and 25.0% for parameters and 33.0% for methods in Python. This suggests that while both models experience hallucination-related errors, Copilot generally performs better, especially in reducing method misprediction errors.
@@
Intent misuses, which occur when the model generates technically correct APIs but fails to fulfill the intended task, also appear frequently. In Python, this misuse type appears in 30.4% of method predictions and 7.4% of parameter predictions for StarCoder. For Copilot, these numbers are lower, at 25.0% for methods and 6.0% for parameters. In Java, intent misuse is present in 27.0% of method misuses and 5.5% of parameter misuses for StarCoder, compared to 22.0% for methods and 4.0% for parameters for Copilot. This indicates that while both models sometimes misunderstand the intended usage of APIs, Copilot shows better performance in both languages.
The Redundancy misuse, where extra or unnecessary parameters or methods are generated, is another notable error. In Python, redundancy misuses account for 14.8% of parameter errors and 5.3% of method errors for StarCoder, compared to 12.0% for parameters and 7.0% for methods for Copilot. In Java, redundant mistakes occur at a rate of 8.2% for parameters and 10.2% for methods for StarCoder, while Copilot has lower rates of 7.0% for parameters and 12.0% for methods. Redundancy errors typically arise when the model includes superfluous API arguments or calls, sometimes due to overfitting on training data patterns or confusion over which parameters are necessary for a given method.
The Missing Item misuse, where expected parameters or method calls are omitted entirely, is a significant issue in Python for both models, appearing in 26.5% of parameter-related errors for StarCoder and 29.0% for Copilot. In Java, missing parameter misuse is less common, with 15.0% for StarCoder and 15.0% for Copilot. This suggests that Python may pose additional challenges to the models when it comes to identifying all required inputs, possibly due to its dynamic typing system or the variation in how functions and methods are structured.
{findingbox}
Finding 2: In general, the most frequent API misuse in LLM-generated code is hallucinations. Hallucination and intent misuses are more prominent in method generation, with Python experiencing higher rates of hallucination errors. Copilot generally performs better than StarCoder, with higher acceptable rates and lower intent and hallucination misuses. Redundancy and missing item misuses are common in both languages, especially for parameters, with Python showing a unique challenge in the form of missing parameter-related misuses.
@@ 5 Mitigation
Mitigating LLM-based API misuse in real-world applications presents several challenges that must be effectively addressed. While automatic misuse detection methods exist, they often necessitate significant manual effort to rectify API misuse. This underscores the need for an automatic approach to aid developers in repairing API misuse. However, the realm of APR methods tailored specifically for API misuse remains largely unexplored. The only relevant work is [13], where the authors evaluated various learning-aid APR models by fine-tuning pre-trained language models on a large-scale (human-written) git commit dataset. However, the study only empirically studied APR for Java and cannot be extended to various programming languages, including Python.
Repairing many LLM-based API misuses requires a nuanced semantic understanding, particularly in deciphering the implicit intents within the code context. Traditional APR methods, reliant on language-specific static analysis or heuristic rules, struggle without delving into these contextual nuances. Moreover, the nondeterministic and cross-lingual nature of libraries utilized in source code poses a formidable challenge for knowledge-constrained APR methods, rendering them ineffective in addressing API misuse across diverse programming environments.
We investigate a promising avenue for API misuse repair - Prompt-based LLM-based APR. Unlike conventional learning-based APR methods trained on task-specific datasets, Prompt-based LLM-based APR is poised to offer superior generalization across a myriad of prompts. Previous research has shown that LLMs can self-repair/debug, enhancing the quality of generated content [47].
@@
Simply prompting LLMs to repair API misuses may not always yield reliable results. LLM-based APR methods can struggle with understanding the nuanced context of the code and the specific intent behind API usage. These models may generate repairs that are syntactically correct but semantically inappropriate, failing to align with the developer’s original intentions. Moreover, LLMs may lack the capability to consistently identify and address the underlying issues causing the misuse. To address these limitations, we propose a novel approach: Mitigating API Misuse via Detect-reason-Fix (Dr.Fix).
Figure 9 overviews our Dr.Fix approach, which integrates detection and reasoning capabilities through multiple stages of LLM prompting. The process begins with the detection stage, where LLMs are prompted to recognize potential API misuse within the code. The LLM is prompted with "Identify any incorrect or unsafe API usages in the following code snippet." Once potential misuses are detected, the taxonomy classification stage follows. Here, the LLM classifies these misuses into predefined categories, such as intent and hallucination misuse. In the reasoning stage, the LLM is prompted to validate the detected misuse with a detailed analysis, considering the broader context and the code’s intended functionality. This process is motivated by self-verification reasoning [48], which has been demonstrated in natural language processing. Finally, in the repair suggestion stage, the LLM generates repair suggestions based on its analysis. These suggestions are crafted to be syntactically correct and semantically meaningful, aligning with the developer’s original intentions.
Incorporating Dr.Fix into the development workflow can significantly reduce the manual effort required to rectify API misuses. By automating both the detection and repair processes, Dr.Fix enhances the reliability and efficiency of software development. This novel approach represents a significant advancement in automated program repair, particularly for LLM-based API misuse, and underscores the potential of combining LLM capabilities with sophisticated reasoning techniques to address complex programming challenges.
@@ 5.2 Backbone Instruction-tuned Models
Llama 3.1 [49] is a collection of multilingual LLMs, a collection of pre-trained and instruction-tuned generative models in 8B, 70B, and 405B sizes released by Meta AI. The Llama 3.1 instruction-tuned text-only models (8B, 70B, 405B) are optimized for multilingual dialogue use cases and outperform many available open-source and closed-chat models on common industry benchmarks. Compared to the original Llama 3 series, Llama-3.1 is significantly improved in terms of math and reasoning capabilities. In our evaluation, we use Llama-3.1 70B, a variant achieving state-of-the-art (SOTA) performance on various benchmarks.
Qwen2.5-Coder
Qwen2.5-Coder [50] is a series of code-specific generative models designed to excel in a wide range of programming tasks such as code generation, completion, reasoning, and repair. Built upon the Qwen2.5 architecture, these models come in sizes ranging from 0.5B to 32B parameters and have been pre-trained on a vast corpus of over 5.5 trillion tokens. Through a combination of extensive data cleaning, scalable synthetic data generation, and balanced data mixing, Qwen2.5-Coder achieves SOTA performance across more than 10 code-related benchmarks. It consistently outperforms other models of similar size, showcasing impressive capabilities in code intelligence while retaining strong general and math skills. With its permissive licensing, Qwen2.5-Coder is positioned to drive innovation in code-related AI research and be widely adopted by developers in real-world applications. We chose Qwen2.5-Coder 32B, the most capable one among the series.

### v2

@@ 4.1 RQ1: How well do LLMs handle API invocation in code generation?
TABLE I: Performance comparison of StarCoder, Copilot, and Qwen2.5-Coder in Python and Java for method and parameter prediction tasks. EM is reported as percentage over 3,000 samples.
@@
BLEU CodeBERTScore EM (%) BLEU CodeBERTScore EM (%)
Method StarCoder 84.6 92.3 75.6 81.2 86.4 44.3
Copilot 88.3 92.6 63.7 86.1 87.7 53.1
Qwen2.5-Coder 92.5 95.2 85.4 91.4 92.3 61.2
Parameters StarCoder 85.0 92.8 77.1 82.0 87.0 44.3
Copilot 89.0 93.0 65.2 87.0 88.2 53.9
Qwen2.5-Coder 93.4 91.1 80.2 93.6 94.7 63.5
@@
We first wanted to investigate how well our studied LLMs can generate code to correctly invoke APIs in code generation. As described above, we sample 3,000 distinct code snippets in Python and Java, respectively, from The Stack dataset. After prompting StarCoder, Copilot, and Qwen2.5-Coder to complete Method Prediction and Parameter Generation independently, we collect 36,000 samples in total. We measure the correctness of the API invocation using three metrics introduced in subsection 3.6: BLEU, CodeBERTScore, and EM. Table V summarizes the performance of the three models on in-the-wild API usage tasks for both Python and Java.
For Python, the results reveal distinct trends across the models. Copilot and Qwen2.5-Coder achieve higher BLEU (88.3 and 92.5) and CodeBERTScore (92.6 and 95.2) than StarCoder (84.6 and 92.3), indicating that they produce more fluent and semantically consistent completions. However, StarCoder maintains competitive exact match performance (75.6%), while Qwen2.5-Coder achieves the highest EM (85.4%), suggesting that it not only generates semantically correct invocations but also matches human-written references more precisely. This improvement likely results from Qwen2.5-Coder’s stronger multilingual code representation and its extensive code-centric training corpus. The divergence between EM and the other metrics highlights how LLMs may generate plausible but syntactically different API invocations, a challenge that remains especially noticeable in Copilot’s outputs.
For Java, both Copilot and Qwen2.5-Coder outperform StarCoder across all metrics. Copilot achieves an EM of 53.1%, while Qwen2.5-Coder achieves 61.2%, compared to StarCoder’s 44.3%. Similar trends are observed for BLEU and CodeBERTScore, where Qwen2.5-Coder leads (BLEU 91.4, CodeBERTScore 92.3), followed by Copilot (86.1 and 87.7) and StarCoder (81.2 and 86.4). This indicates that Qwen2.5-Coder, although not instruction-tuned, generalizes better to Java APIs, possibly due to broader coverage and better pretraining on strongly typed languages.
The comparison among these three decoder-based models highlights how pretraining strategies and dataset diversity shape model behavior. StarCoder shows strong memorization and syntactic alignment in Python due to its Python-focused fine-tuning. Copilot achieves consistent semantic fluency across languages through large-scale, multi-language exposure, while Qwen2.5-Coder balances both, delivering strong exact matches and high fluency simultaneously. This suggests that newer open-source models are closing the performance gap with proprietary systems in real-world API invocation tasks.
StarCoderCopilotQwen2.5-Coder126466152646481278
(a) Wrong Method Infilling in Python
StarCoderCopilotQwen2.5-Coder1384305417298163279
(b) Wrong Parameter Generation in Python
StarCoderCopilotQwen2.5-Coder2332472514203148721
(c) Wrong Method Infilling in Java
StarCoderCopilotQwen2.5-Coder443232109388223146617
(d) Wrong Parameter Generation in Java
@@
Figure 8: Overlap of incorrect API usages among StarCoder, Copilot, and Qwen2.5-Coder for method infilling and parameter generation in Python and Java.
Figure 8 shows the overlap and discrepancies in API misuse among the three models. In Python Method Infilling (8a), StarCoder has 126 unique errors, Copilot 466, and Qwen2.5-Coder 15, with substantial overlaps where 278 errors are shared by all three models. Similar patterns hold for Java Method Infilling (8c), where StarCoder has 233 unique errors, Copilot 24, and Qwen2.5-Coder 72, again with a large shared subset (721). For parameter generation, the models exhibit analogous overlap patterns where shared errors dominate, although Qwen2.5-Coder consistently produces fewer unique misuses. These results indicate that while Qwen2.5-Coder shows fewer unique errors, all three models converge on many of the same failure cases, reflecting persistent limitations in LLM reasoning about API semantics. The observed differences can be traced back to each model’s training focus. StarCoder’s specialization in Python contributes to high syntactic fidelity but limited cross-language transfer, Copilot benefits from multi-language fine-tuning but sacrifices exactness, and Qwen2.5-Coder attains strong generalization through diverse code-centric pretraining without instruction tuning.
@@
Finding 1: Qwen2.5-Coder achieves the highest overall accuracy across both languages, combining strong fluency with precise matching. Copilot follows closely, while StarCoder remains competitive in Python due to domain specialization. Despite these differences, all three models share a substantial number of common API misuses, which shows that correct API invocation remains a fundamental challenge for current LLMs.
@@ 4.2 RQ2: How do LLMs misuse APIs in the wild?
TABLE II: Comparison of API Misuse by StarCoder in Java and Python. “None” indicates that the LLM-based API usage was considered acceptable based on human inspection.
@@
TABLE III: Comparison of API Misuse by Copilot in Java and Python. “None” indicates that the LLM-based API usage was considered acceptable based on human inspection.
@@
None 35.0 24.5 28.0 27.3
TABLE IV: Comparison of API Misuse by Qwen2.5-Coder in Java and Python. “None” indicates that the LLM-based API usage was considered acceptable based on human inspection.
Misuse Type Java Python
Params (%) Methods (%) Params (%) Methods (%)
Intent 4.2 20.5 6.5 23.6
Hallucination 33.5 26.8 23.0 29.4
Redundancy 6.3 10.5 10.7 6.4
Missing Item 13.0 11.7 25.4 8.1
None 43.0 30.5 34.4 32.5
We further analyze the patterns and distributions of API misuse across StarCoder, Copilot, and Qwen2.5-Coder. Specifically, we study the annotated outputs introduced in subsection 3.7, which correspond to samples that failed on EM score. Table II, Table III, and Table IV show the distribution of API misuse types across Java and Python, categorized by methods and parameters.
Hallucination remains the most frequent misuse across all three models, although its prevalence varies. For StarCoder, hallucinations dominate all configurations, accounting for up to 43.3% of parameter misuses in Java. Copilot shows slightly reduced rates, while Qwen2.5-Coder exhibits further decreases, with hallucination rates below 30% in most cases. This indicates that Qwen2.5-Coder is better grounded in valid API namespaces, reducing the tendency to generate non-existent or version-inconsistent APIs.
Missing Item misuses are prominent for Python parameters, particularly for Copilot and Qwen2.5-Coder. Copilot omits required parameters in 29.0% of Python parameter cases, making it the leading error type for that configuration. Qwen2.5-Coder also shows relatively high omission rates (25.4%), although slightly lower than Copilot’s. This pattern suggests that the challenge of inferring correct argument lists persists even for newer models, especially in Python where argument flexibility and version heterogeneity increase complexity.
Intent misuses continue to appear frequently but are less severe in Qwen2.5-Coder. In Python, StarCoder records 30.4% intent errors for methods, while Copilot and Qwen2.5-Coder reduce these to 25.0% and 23.6% respectively. Similar trends appear in Java. These improvements suggest that newer models capture functional intent more reliably, although confusion between similar API methods still occurs.
Redundancy misuses are less frequent but remain consistent across models. Qwen2.5-Coder reduces redundant arguments and superfluous method calls compared to StarCoder and Copilot, showing the lowest rates overall (around 6–10%). This aligns with its improved parameter consistency and contextual inference capabilities.
Overall, Python parameter completion remains the most error-prone scenario. The results across all three models indicate that Python’s dynamic typing and overlapping library interfaces still pose major challenges. Frequent hallucination and omission misuses demonstrate that LLMs struggle to fully align API invocations with real-world library constraints, especially for parameters that require contextual understanding of object states or external imports.
{findingbox}
Finding 2: Hallucination remains the most common misuse overall, although Qwen2.5-Coder reduces its occurrence compared with StarCoder and Copilot. Missing Item errors remain prominent in Python parameter prediction for all models. Intent misuses are frequent in method prediction, but Qwen2.5-Coder demonstrates improved grounding and fewer redundant errors, suggesting gradual progress toward more reliable API usage.
@@ 5 Mitigation
Mitigating LLM-based API misuse in real-world applications presents several challenges. Existing automatic misuse detection methods often require substantial manual effort to apply fixes, underscoring the need for approaches that can both detect and repair misuses. However, program repair methods tailored specifically for API misuse remain largely unexplored. The only relevant work is [zhang2023evaluating], which evaluated pre-trained models fine-tuned on large-scale git commit patches, but focused only on Java. More recently, Wei et al. proposed LLMAPIDet [wei2024demystifying], a detector and repair tool for deep learning libraries based on ChatGPT. While effective in that domain, it is limited by its reliance on closed-source models and its focus on TensorFlow and PyTorch APIs. By contrast, Dr.Fix is designed to be backbone-agnostic and taxonomy-driven, making it applicable to general-purpose third-party APIs across different programming languages.
Repairing LLM-generated API misuses is particularly important because such misuses often arise from the way LLMs generate code, rather than from human misunderstanding of documentation. Unlike traditional APR methods, which rely on static analysis or heuristics, LLM-based repair can leverage contextual knowledge of APIs and natural language reasoning to propose fixes. This enables handling of misuse types such as hallucinations or incorrect intent selections, which are less common in human-written code but prevalent in LLM outputs. At the same time, it is valuable to investigate whether LLMs reproduce the same kinds of misuses as humans, or if they introduce distinctive patterns that call for new repair strategies.
We explore a promising direction for API misuse repair: prompt-based LLM program repair. Unlike conventional APR methods trained on task-specific datasets, prompt-based approaches offer greater flexibility and generalization. Prior work has shown that LLMs are capable of self-repair and debugging through structured prompting [chenteaching], motivating our design of Dr.Fix.
@@
Repairing LLM-based API misuses requires nuanced semantic understanding, particularly of the implicit intents within the code context. While LLMs can generate repairs when directly prompted, prior work shows that naive prompting often produces syntactically valid but semantically incomplete or inconsistent fixes [chenteaching, sclarquantifying, miaoselfcheck, mizrahi2024state]. Moreover, recent studies on domain-specific APIs highlight that LLM-based repair is sensitive to how misuses are detected and framed, and simple prompts may overlook subtle but critical constraints [wei2024demystifying, kwon2023exploring]. These observations motivate the need for structured prompting strategies that incorporate detection, classification, and reasoning, rather than relying on single-pass repair. At the same time, traditional APR methods, reliant on static analysis or heuristic rules, remain limited when applied across diverse programming environments.
Figure 9 overviews our Dr.Fix approach, which integrates detection and reasoning capabilities through multiple stages of LLM prompting. The process begins with the detection stage, where the model is prompted to recognize potential API misuse within the code. Once potential misuses are identified, a taxonomy classification stage follows, mapping them into predefined categories such as intent or hallucination misuse. In the reasoning stage, the model validates the detected misuse with a detailed analysis that considers the broader context and intended functionality. This stage is motivated by self-verification reasoning [weng2023large], which has been effective in natural language processing tasks. Finally, the repair suggestion stage produces candidate fixes that aim to be both syntactically correct and semantically meaningful, aligning with the developer’s intent.
Incorporating Dr.Fix into the development workflow reduces the manual effort required to address API misuses. By automating both detection and repair, Dr.Fix improves the reliability and efficiency of software development. Our design demonstrates the potential of combining LLM capabilities with structured reasoning techniques to address complex challenges in automated program repair.
@@ 5.2 Backbone Instruction-tuned Models
LLMs serve distinct purposes in software engineering workflows. Models designed for IDE-based code completion prioritize efficiency and responsiveness, enabling seamless interaction within development environments. In contrast, models aimed at code repair focus on semantic correctness and reliability rather than generation speed, as repair involves complex reasoning over code context and intent. Table V summarizes the speed and cost characteristics of representative models in both categories. While the IDE-based code completion results are derived from our previous experiments, in this work we concentrate on LLMs for API misuse repair when evaluating the performance of Dr.Fix. To ensure the highest possible repair quality, we adopt frontier LLMs which better represent the current technological upper bound in reasoning and code understanding. This distinction allows us to balance efficiency and precision while analyzing the contribution of repair-oriented LLMs to Dr.Fix’s overall effectiveness.
TABLE V: Model Comparison: Speed and Cost. Speed and cost for all models except StarCoder and Copilot are obtained from OpenRouter [openrouter]. For StarCoder and Copilot, we measure inference speed locally; both are free to use in this setup. We note that the cost is measured per 1M tokens.
IDE
Metric StarCoder Copilot Qwen2.5-Coder
Speed (tokens/s) 217.4 135.0 229.0
Cost (\$ in/out) 0.0 / 0.0 0.0 / 0.0 0.03 / 0.09
Repair
Metric Llama 3.1 Qwen2.5-Coder-Instruct GPT-4o
Speed (tokens/s) 22.65 38.17 59.3
Cost (\$ in/out) 0.4 / 0.4 0.04 / 0.16 2.5 / 10
@@
Llama 3.1 [dubey2024llama] is a collection of multilingual LLMs, a collection of pre-trained and instruction-tuned generative models in 8B, 70B, and 405B sizes released by Meta AI. The Llama 3.1 instruction-tuned text-only models (8B, 70B, 405B) are optimized for multilingual dialogue use cases and outperform many available open-source and closed-chat models on common industry benchmarks. Compared to the original Llama 3 series, Llama-3.1 is significantly improved in terms of math and reasoning capabilities. In our evaluation, we use Llama-3.1 70B, a variant achieving state-of-the-art (SOTA) performance on various benchmarks.
Qwen2.5-Coder-Instruct
Qwen2.5-Coder-Instruct [hui2024qwen2] is a series of code-specific generative models designed to excel in a wide range of programming tasks such as code generation, completion, reasoning, and repair. Built upon the Qwen2.5 architecture, these models come in sizes ranging from 0.5B to 32B parameters and have been pre-trained on a vast corpus of over 5.5 trillion tokens. Through a combination of extensive data cleaning, scalable synthetic data generation, and balanced data mixing, Qwen2.5-Coder-Instruct achieves SOTA performance across more than 10 code-related benchmarks. It consistently outperforms other models of similar size, showcasing impressive capabilities in code intelligence while retaining strong general and math skills. With its permissive licensing, Qwen2.5-Coder-Instruct is positioned to drive innovation in code-related AI research and be widely adopted by developers in real-world applications. We chose Qwen2.5-Coder-Instruct 32B, the most capable one among the series.